OpenStudy (anonymous):

W+4 3 -------- = -- 4W+6 4

OpenStudy (janu16):

is 3 and 4 a fraction?

OpenStudy (kittiwitti1):

I think she meant to type \[\frac{W+4}{4W+6}=\frac{3}{4}\]

OpenStudy (janu16):

is tht waht you meant?

OpenStudy (anonymous):

Yah that's what I meant

OpenStudy (janu16):

you solving for w? right?

OpenStudy (anonymous):

How did you make it look like that?

OpenStudy (anonymous):

And yes I am

OpenStudy (janu16):

theres a blue button that says equation so click tht and you can make it look likwe tht

OpenStudy (anonymous):

I guess they don't have it for the iPhone

OpenStudy (kittiwitti1):

I actually typed out LaTex, which is the font format for all that junk lol I think they have a little training forum thread for that. ^-^

OpenStudy (janu16):

ok so you need help right?

OpenStudy (anonymous):

Yah I'm not sure how to go about it if you could just help get it to a regular problem I could figure it out

OpenStudy (kittiwitti1):

multiply both sides by the denominator of the left, then combine like terms and isolate W

OpenStudy (anonymous):

Hats after that?

OpenStudy (kittiwitti1):

\[(4W+6)\left(\frac{W+4}{4W+6}\right)=(4W+6)\left(\frac{3}{4}\right)\]

OpenStudy (anonymous):

*whats

OpenStudy (kittiwitti1):

Left side cancels, so we get W+4.\[W+4=\frac{3(4W+6)}{4}\]

OpenStudy (kittiwitti1):

Can you do it from here? Just distribute 3 there...

OpenStudy (anonymous):

Yah I can thank you guys for the help!